OpenStudy (anonymous):

If h(x) = −4x − 10, find h(−5). −30 1.25 10 −1 PLease help I will give a medal

OpenStudy (bradely):

Substitute x = -5 in -4x - 10 -4(-5) -10 = 20-10 =10 Any help post on mathskey.com